Consider the following equations:

SUPPLY: Q=10+2P

DEMAND: Q=60-3P

a) Compute the consumer's surplus at equilibrium. Show it on a graph.
Answer #1

Setting QD=QS

60-3P=10+2P

60-10=2P+3P

50=5P

P=10

Q=10+2(10)=30

When QD=0,P=60/3=20

CS=0.5X30X(20-10) = 150

CGE tO 30